The density of Kcl is 1.9893 gm/cm3 .If the atomic mass of K and Cl are 39 amu and 35.5 amu respectively, find out the inter ionic distance of KCl.

Grade:12

1 Answers

Chhavindra Tripathi
5 years ago
d=Z*M/ NA*a3……..(1)
density d= 1.98 gm/cm3
Molar mass M= 39+35.5=74.5
Structure of KCl is NaCl ype so Z=4
Put in equation 1 and find a
a= 6.28*10^-8 cm
inter ionic distance in NaCl r++r--= a/2 = 3.14*10^-8 cm